OpenStudy (anonymous):

SOMEONE PLZ HELP!!! what is the simplified form of 4x + 2 over x + 5 - 3x - 1 over x + 5 ???

OpenStudy (anonymous):

first combine like terms and get 4x+2 over 4x+4 over x+5

OpenStudy (anonymous):

\[\frac{4 x+2}{x+5}-\frac{3 x-1}{x+5}=\frac{-(3 x-1)+4 x+2}{x+5} \]

OpenStudy (anonymous):

\[\frac{x+3}{x+5} \]

OpenStudy (anonymous):

Ok, so how did you cancel the numerators to x + 3?